40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9584     Entry time: Tue Jan 28 23:32:12 2014
Author: Koji 
Type: Update 
Category: General 
Subject: X/Y arm locked with the IR beam 

[Koji EricQ]

The both arms have been locked with IR and aligned by ASS.

The IFO was left with ITMX/Y, ETMX/Y, BS, and PRM aligned, and the PSL shutter closed.


YARM
SIGNAL PATH:
POY11I(+45dB)->YARM(G=+1.0)->ETMY
NORM: TRYx10
TRIG: TRY 0.01up/0.005down
FM TRIG: FM2/3/6/7/8/9 0.01up/0.05down, 0.5 sec delay

XARM
SIGNAL PATH:
POX11I(+45dB)->XARM(G=+4.0)->ETMX
NORM: TRXx10
TRIG: TRX 0.01up/0.005down
FM TRIG: FM2/3/6/7/8/9 0.01up/0.05down, 0.5 sec delay


For decent locks, it was necessary that the offset of the error signals are trimmed at the input filters
even after running LSCoffset.py script.

Once the cavities were aligned for the IR, we could see the green beams are also flashing.
The Y arm was actually locked with the green with a TEM00 mode

Attachment 1: good_alignment.png  18 kB  | Hide | Hide all
good_alignment.png
ELOG V3.1.3-